OpenStudy (anonymous):

Use the data in the following table, which summarizes results from 171 pedestrian deaths that were caused by accidents. If two different deaths are randomly select the replacement, find the probability they both involved intoxicated drivers. Pedestrian Intoxicated? Yes No Drivers yes 24 34 Intoxicated? No 97 15 I wonder if I have the right formula!!! Please help me, Thanks!!! ;))

OpenStudy (anonymous):

What do you have so far?

OpenStudy (anonymous):

P(A and B)=P(A) +P(B)-P(A and B)

OpenStudy (anonymous):

Recursive equation?

OpenStudy (anonymous):

122/171 + 59/171-34/171

OpenStudy (anonymous):

P(A or B)?

OpenStudy (anonymous):

By the way, have you learned about mutual exclusiveness and independence?

OpenStudy (anonymous):

Ok. How I used it, when P(A or B)?? No, I might have but, I don't remember.

OpenStudy (anonymous):

So P(A | B) = P(A) + P(B) - P(A & B) give us P(A & B) = P(A) + P(B) - P(A | B). That's what you meant right?

OpenStudy (anonymous):

| for or & for and

OpenStudy (anonymous):

Yes!

OpenStudy (anonymous):

That will give us the probability for one case.

OpenStudy (anonymous):

(By the way, should those number be adding up to 171 or 170)?

OpenStudy (anonymous):

Why 170?

OpenStudy (anonymous):

24+34+97+15 = 170 also 24+34 = 58

OpenStudy (anonymous):

Oohhh!!

OpenStudy (anonymous):

Sorry is actually 25 instead of 24.

OpenStudy (anonymous):

"If two different deaths are randomly select the replacement," Does this mean after we select the death, we can possibly select the same death again?

OpenStudy (anonymous):

Ok

OpenStudy (anonymous):

It's a question, not a statement.

OpenStudy (anonymous):

Ooh! 171

OpenStudy (anonymous):

I am trying to solved it

OpenStudy (anonymous):

Noo, it's the answer.

OpenStudy (anonymous):

The answer is 0.1177 I confuse now..

OpenStudy (anonymous):

So 11.77% chance, right?

OpenStudy (anonymous):

Do you have anyway to verify the answer?

OpenStudy (anonymous):

Nop!

OpenStudy (anonymous):

It was a quiz.

OpenStudy (anonymous):

So, How you got that? I still don't get it. even though I can do the other one, but this one is tricky.

OpenStudy (anonymous):

@rocal2 I think I messed up on it.

OpenStudy (anonymous):

The problem is that I don't understand the exact text of the problem.

OpenStudy (anonymous):

To be honest, me neither. But let me double check if I messed up in writing.

OpenStudy (anonymous):

I'll double check and I didn't messed up a world.

OpenStudy (anonymous):

Sorry but the English is a bit bad on the question. That is why my problem is.

OpenStudy (anonymous):

I think the answer is [(25+34)/171]^2

OpenStudy (anonymous):

But it could be wrong because the question is unclear.

OpenStudy (anonymous):

Don"t worry, I have the same situation with English as second language..Is difficult for me too.

OpenStudy (anonymous):

Still, not that one..

OpenStudy (anonymous):

How about (25+34)/171 * [(25+33)/170 + (24+34)/170]

OpenStudy (anonymous):

Nopi!!

OpenStudy (anonymous):

I can't do it. I simply don't know what it is asking. Sorry.

OpenStudy (anonymous):

It's ok! don't worry. Thanks for trying anyway.. ;))